Pure and dry nitrogen gas is slowly bubbled through a solution of 2.25 g of a non- volatile organic compound dissolved in 150g of benzene and then bubbled through pure benzene. The solution is found to be 2.1540g lighter, while pure benzene suffered a loss in weight of 0.0160 g. Calculate the molecular weight of non-volatile organic compound.


1
Expert's answer
2021-01-26T03:10:56-0500

2.25/x – moles of the compound, where x – molar mass of it.

150/78.11 = 1.92 – moles of benzene, where 78.11 – molar mass of it.

 

– mole fraction of the solution

The comparison of two weight losses gives: 0.0160 g / 2.1540 g = 0.007428

 


x = 156.6 (g/mol)